In a certain Algebra 2 class of 29 students, 7 of them play basketball and 14 of them play baseball. There are 10 students who play neither sport. What is the probability that a student chosen randomly from the class plays basketball or baseball?



1
Expert's answer
2021-02-01T08:05:14-0500

Number of students = 29

7 of them play basketball and 14 of them play baseball. There are 10 students who play neither sport.

Number of students playing = 29 – 10 = 19

Number of students playing both games = 7 + 14 – 19 = 2

Number of students playing only basketball = 7 – 2 = 5

Number of students playing only baseball = 14 – 2 = 12

Number of students playing basketball or baseball = 5 + 12 = 17

Probability that a student chosen randomly from the class plays basketball or baseball:

"P = \frac{17}{29} = 0.586"

Answer: 0.586